OpenStudy (anonymous):

Completely factor the following expression: 49y²– 81z²

jimthompson5910 (jim_thompson5910):

49y^2-81z^2 (7y)^2-81z^2 (7y)^2-(9z)^2 Now use the difference of squares rule A^2-B^2=(A+B)(A-B) to factor. In this case, A = 7y and B = 9z